I need to use the change of base formula to solve 3^2x=15
Answer by Alan3354(21583)

3^2x=15
---------
To solve for x:
log(3^2x) = log(15) common logs, base 10. Natural logs would work too.
2x*log(3) = log(15)
2x = log(15)/log(3)
x = log(15)/2log(3)
x = log(15)/log(9)
x = ~ 1.2325